Grade 12th passAlgebra

√x+√x+√x+√x+.................∞ = 6 find the value of x

ER Munna Muskan
9 Years agoGrade 12th pass

1 Answer

Arun
8 Years ago
because there are infinite terms
\sqrt{}x+\sqrt{}x+\sqrt{}x+.........infinite terms = 6
\sqrt{}(x+6) = 6
x+6 = 36
x =30
Note- This is a simple approach to this type of all questions.